This has me baffled. R=PV/nT;V (Chemistry)
Options: A) V=n/PRT
B) V=nRT/P
C) V=nP/RT
D) V=n/P
E) V=P/nRT

Answer by user_dude2008(1862)

R=PV/nT

RnT=PV

V=(RnT)/P

V=nRT/P